OpenStudy (anonymous):

When evaluated, which expression has a result that is a rational number? a. √3 + 6 b. √36 - π c. √144 x 0.25 d. √4 ÷ √2

OpenStudy (michele_laino):

hint: \(\pi\) is an irrational number, furthermore also \( \sqrt 3 \) and \( \sqrt 2\) are irrational numbers. Finally: \(\sqrt 144=12\) and \( \sqrt 4 = 2\)